40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 16 16:00:01 2019, gautam, Update, CDS, c1iscaux Supermicro setup IMG_7769.JPG
    Reply  Wed Jul 17 21:22:41 2019, gautam, Update, CDS, CM board Latch Enable subtlety 
       Reply  Sun Jul 21 12:55:38 2019, gautam, Update, CDS, CM board Latch Enable test script LatchLogic.pdfLatchLogicTest.png
          Reply  Fri Aug 9 08:59:04 2019, gautam, Update, CDS, Prep for install of c1iscaux 
             Reply  Sun Aug 11 11:47:42 2019, gautam, Update, CDS, Bench test of c1iscaux 
                Reply  Mon Aug 12 17:36:04 2019, gautam, Update, CDS, More bench test of c1iscaux iscauxCheclist.pdf
                   Reply  Mon Aug 12 21:25:19 2019, Koji, Update, CDS, More bench test of c1iscaux 
                   Reply  Tue Aug 13 08:07:09 2019, gautam, Update, CDS, P1--->P2 
                      Reply  Tue Aug 13 14:36:17 2019, gautam, Update, CDS, P1--->P2 
                   Reply  Fri Aug 16 16:40:04 2019, gautam, Update, CDS, 1Y3 work newLook1Y3.JPGIMG_7803.JPGc1lsc_crashed.png
                      Reply  Sat Aug 17 16:49:23 2019, gautam, Update, CDS, More 1Y3 work Screen_Shot_2019-08-17_at_3.00.57_PM.pngScreen_Shot_2019-08-17_at_3.12.23_PM.pngIMG_7804.JPGScreenshot_from_2019-08-17_17-04-47.png
                         Reply  Mon Aug 19 14:36:21 2019, gautam, Update, CDS, c1iscaux remaining work caseForSmallerFootprint.pdf
                            Reply  Tue Aug 20 19:05:24 2019, Koji, Update, CDS, MC1 (and MC3) troubleshoot Screenshot_from_2019-08-20_17-26-01.pngScreenshot_from_2019-08-20_17-43-03.png
                               Reply  Thu Aug 22 12:54:06 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 7x
                                  Reply  Thu Aug 22 20:56:51 2019, Koji, Update, CDS, MC1 glitch removed (for now) and IMC locking recovered 
                            Reply  Fri Aug 23 18:46:17 2019, Jon, Update, CDS, c1iscaux remaining work 
                               Reply  Sun Aug 25 14:18:08 2019, gautam, Update, CDS, c1iscaux remaining work Screen_Shot_2019-08-25_at_10.38.37_PM.png
    Reply  Thu Jul 18 10:46:04 2019, gautam, Update, CDS, Database files made 
       Reply  Thu Jul 18 19:58:56 2019, gautam, Update, CDS, Work on Acromag chassis IMG_7771.JPGIMG_7770.JPGIMG_7772.JPG
          Reply  Fri Jul 19 19:44:03 2019, gautam, Update, CDS, Database file test Whitening.png
Message ID: 14840     Entry time: Sun Aug 11 11:47:42 2019     In reply to: 14837     Reply to this: 14841
Author: gautam 
Type: Update 
Category: CDS 
Subject: Bench test of c1iscaux 

I bench tested the functionality of all the c1iscaux Acromag crate channels. Summary: we are not ready for a Monday install, much debugging remains.

  1. DAC channels were tested using 4 ch oscilloscope and stepping the whitening gain sliders through their 15 gain settings
    • Response was satisfactory - the output changes between 0 - 5 V DC in 15 steps.
    • This analog voltage is converted to binary representation by an on-board ADC on the whitening boards. So we may have to tune the offset voltage and range to avoid accidental bit flipping due to the analog voltage of a particualr step falling close to the bit-flipping edge of the on-board ADC. This will require an in-situ test.
    • Test passed
  2. BIO output channels were tested using a DMM, and monitoring the resistance between the BIO pin and the RTN pin. In the "ON" state, the expected resistance is ~5 Mohm, and in the off state, it is ~3 ohms.
    • The AA filter switches on BIO1 unit do not show the expected behavior - @ Chub, please check the wiring.
    • All others (except the mbboDirect bits, see next bullet) were okay, including those for the CM board that are NOT part of the mbboDirect groups.
    • Test failed
  3. ADC channels were tested by driving a ~2Vpp 300mHz sine wave with a function generator, and looking at the corresponding EPICS channel with StripTool.
    • I found that all the ADC channels don't function as expected.
    • Part of the problem is due to incorrect formatting of the EPICS records in the db files, but I think the ADCs also need to be calibrated with the precision voltage source.
    • Why only ADCs require calibration and not the DACs????
    • Test failed
  4. mbboDirect BIO output test - I made a little LED breadboard tester kit to simultaneously monitor the status of these groups of binary outputs.
    • The LSB is toggled as expected when moving the gain slider along.
    • However, the other bits in the group are not toggled correctly.
    • I believe this is a problem with either (i) the way the EPICS record is configured to address the bits or (ii) the incorrect modbus datatype is used to initialize the ioc.
    • It will be helpful if someone can look into this and get the mbboDirect bits working, I don't really want to spend more time on this.
    • Test failed

I am leaving the crate powered (by bench supplies) in the office area so I have the option to work remotely on this.

ELOG V3.1.3-